### v1

# No Train but Gain: Language Arithmetic for training-free Language Adapters enhancement

Abstract
Modular deep learning is the state-of-the-art solution for lifting the curse of multilinguality, preventing the impact of negative interference and enabling cross-lingual performance in Multilingual Pre-trained Language Models. However, a trade-off of this approach is the reduction in positive transfer learning from closely related languages. In response, we introduce a novel method called language arithmetic, which enables training-free post-processing to address this limitation. Inspired by the task arithmetic framework, we apply learning via addition to the language adapters, transitioning the framework from a multi-task to a multilingual setup. The effectiveness of the proposed solution is demonstrated on three downstream tasks in a MAD-X-based set of cross-lingual schemes, acting as a post-processing procedure. Language arithmetic consistently improves the baselines with significant gains in the most challenging cases of zero-shot and low-resource applications. Our code and models are available at https://github.com/mklimasz/language-arithmetic.
1 Introduction
The recent progress of large language models has led to a huge interest in how well they can perform not just in English but across multiple languages and invoked the rise of Multilingual Pre-trained Language Models (MLLMs) [11, 43, 2]. These models serve as general-purpose solutions that can be adapted and applied to various Natural Language Processing tasks. Notably, MLLMs demonstrate zero-shot cross-lingual capabilities, allowing them to generalise effectively to downstream tasks even when trained in a language different from the target language.
The positive transfer of abilities from both related languages and high-quality training data from unrelated languages has meant that MLLMs have reported state-of-the-art performance in low-resourced languages [24]. However, this benefit does not always extend to high-resourced languages [21]. In such cases, the quality of MLLMs tends to decrease compared to their monolingual counterparts [25, 23, among others] due to negative interference phenomena [42]. Additionally, the curse of multilinguality [11] reveals the existence of a trade-off between language coverage and model capacity. Consequently, MLLMs must carefully limit the number of languages included during the pre-training phase.
Modular deep learning (MDL) [32] methods come to the rescue to avoid the abovementioned limitations, enabling the extension of MLLMs to support any number of languages [6, 40, 34, 29, 31]. MDL methods adapt the model to arbitrary tasks and languages by isolating components from each other (and the backbone MLLM) via parameter-efficient extensions. Examples of parameter-efficient modules are LoRA [17] and adapters [37, 16], which are low-budget (in terms of parameters) bottleneck layers that increase an MLLM size by just a fraction. Language adapters especially allow the modularisation of language-specific knowledge by training on a raw, unlabelled corpus in a target language.
The limitation of the MDL and language adapters is their isolation. While they lift the curse of multinguality and prevent negative interference, at the same time, language adapters limit the possible impact of positive transfer. Previous attempts to address these challenges — such as training bilingual [26] or language-family [9] adapters — do not scale effectively. In our work, we tackle this limitation as a post-processing step. Leveraging recent insights from task arithmetic [19], specifically learning via addition, we augment language adapters with missing related language knowledge – a concept we term language arithmetic. Remarkably, this training-free approach enhances language adapters across various scenarios, from zero-shot settings to low-resource regimes.
To summarise, our contributions are as follows:
  • A novel training-free post-processing method named language arithmetic that enhances language adapters.
  • We conduct a cross-lingual evaluation on three downstream tasks and Multilingual Pre-trained Language Models with test cases that include zero-shot and low-resource setups.
  • We provide an analysis of language arithmetic internal components.
2 Background
Our research builds upon the contributions of Ilharco et al. [19] and Zhang et al. [45]. The following Section provides the background and serves as a gentle introduction to the concept of task vectors and task arithmetic.
2.1 Task Vectors & Task Arithmetic
Let us assume that we have access to a pre-trained model with its weights denoted $\theta_{pre}\in R^{d}$ and a fine-tuned version of the same model on a task $t$ represented by $\theta^{t}_{ft}\in R^{d}$. The task vector $\tau_{t}\in R^{d}$ is an element-wise difference between models’ weights.
$\tau_{t}=\theta^{t}_{ft}-\theta_{pre}$ (1)
The task vectors can be part of multiple arithmetic operations, e.g. learning via addition. This operation is an addition operation between two task vectors and the base model, i.e. we add two differences between the fine-tuned models and the pre-trained version with weights controlling the impact.
$\theta_{multi-task}=\theta_{pre}+\lambda_{1}\tau_{t_{1}}+\lambda_{2}\tau_{t_{2}}$ (2)
The lambdas can be further normalised to sum to one, i.e. $\lambda_{2}=1-\lambda_{1}$ and simplifying notation with just $\lambda$.
$\theta_{multi-task}=\theta_{pre}+\lambda\tau_{t_{1}}+(1-\lambda)\tau_{t_{2}}$ (3)
While we define learning via addition for two tasks, the same procedure can be applied to multiple tasks.
Task arithmetic allows us to forge a multi-task model from a separate, task-specific set of fine-tuned models, preserving high accuracy (although a shared pre-trained starting point is required, e.g. the same Language Model). Moreover, vectors from different tasks are typically close to orthogonal, and Ilharco et al. [19] speculate that this enables the combination of task vectors via addition with minimal interference.
In the Parameter-Efficient Fine-tuning (PEFT) regime, we can reduce a task vector to only newly introduced PEFT parameters [45]. Another way of looking at this is a task vector that is a zero-valued vector for all the parameters apart from PEFT weights, i.e. a task vector would not modify the base, pre-trained checkpoint.
3 Method
We propose language arithmetic that transitions the task arithmetic concept from a multi-task to a multilingual setup. In this Section, we describe the language arithmetic alongside its application as a training-free, post-processing step to a MAD-X cross-lingual framework [29].
3.1 Language Arithmetic
We formulate a language arithmetic (LA) concept by substituting the task in task vectors and arithmetic with a language. This approach means that instead of merging downstream tasks, we target a problem of cross-lingual performance. We propose to apply learning via addition to languages, and in Section 5.3, we demonstrate the discrepancies when comparing language and task vectors. Our study focuses specifically on the language adapters [34, 29]. Due to overlapping abbreviations, we use the LA exclusively as the former, i.e. language arithmetic. In the learning via addition, we limit the parameters to language adapters and simplify the notation that $\theta$ represents the adapters’ weights and $\tau$ is referred to as a language vector. As we operate in a language space instead of a task, the $t$ is replaced with a language, i.e. its language code in the notation. The example equation describes a language arithmetic operation between an English and a Spanish adapter.
$\theta_{LA}=\theta_{pre}+\lambda\tau_{en}+(1-\lambda)\tau_{es}$ (4)
Throughout the paper, the equation above is abbreviated as a function: $LA(en,es)$ with lambda as a default parameter. Whenever we use language arithmetic in a zero-shot context, we add a subscript to the function name, i.e. $LA_{0s}(en,es)$ means that language arithmetic between English and Spanish is applied to a different language, e.g. French.
Language arithmetic is a training-free method, taking advantage of already pre-trained modules. The sole requirement is a validation dataset on which the $\lambda$ parameter can be established. While in our work, we use a pretty fine-grained step ($0.05$) to determine the $\lambda$ value (i.e. we run evaluations for $\lambda\in[0,1]$ with a provided step), our analysis showcased that it is possible to increase the value and limit the computation burden even more (details in Section 5.1).
3.2 Application
We apply the language arithmetic as an enhancement method for language adapters. We assume two scenarios: (i) zero-shot case, where we do not have a language adapter in the desired language and (ii) “regular” enhancement case, where we improve the existing language adapter (in high- and low-resourced languages).
We evaluate our post-processing method as an extension of the MAD-X framework [29] to challenge our solution in a cross-lingual manner. The overview of the schema is presented in Figure 1.
[…]
The MAD-X consists of the following:
  1. 1.
    Choosing MLLM
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
[…]
In our work, we introduce an additional step:
  1. 5.
[…]
3.2.1 MLLM
The first step usually assumes the availability of a Multilingual Pre-trained Language Model. In our work, we focus on mBERT111bert-base-multilingual-cased [12] and XLM-R222xlm-roberta-base [11].
3.2.2 Training language & task adapters
The following two steps depend on the downstream dataset. In the most extreme case, we might have a dataset in one specific language and need to take advantage of cross-lingual performance. Therefore, step number 2 trains a language adapter on raw corpora corresponding to the task language using regular masked language modelling loss in a self-supervised manner. During this step, the MLLM is frozen. The third step freezes both the backbone and language adapter and trains a task adapter on a downstream task dataset. Given a set of tasks or if a new task appears, we can repeat this step as long as a required language adapter exists, i.e. a language adapter that matches the task’s language.
3.2.3 Additional language adapters
The last, fourth step is the heart of the MAD-X framework. We can now train language adapters in any language, leverage the pre-trained task adapter(s) and obtain a cross-lingual performance by connecting a new language adapter with a task adapter (i.e. routing first via language adapter and then task) to obtain language-task pair unseen during training. The growing pool of pre-trained adapters can be accessed at public repositories like AdapterHub [28] and reused for further use cases.
3.2.4 Post-processing via Language Arithmetic
First, we assume the use-case where the pool of language adapters does not contain the target languages, i.e. zero-shot scenario. In this study, we add English as a task-specific language to a language related to the target. In the second language adapter enhancement scenario, we apply language arithmetic in its default form, where the goal is to enhance the existing target language adapter (via either English or a related language). At last, we investigate (simulated and actual) low-resource cases.
4 Experiments
4.1 Experimental setup
We use the AdapterHub library [28] for all our experiments. For language adapters, we train on the Wikipedia corpora33320331101.xx checkpoint https://huggingface.co/datasets/wikimedia/wikipedia for 250k steps with a learning rate of 1e-4, an effective batch size of 64 using a single GPU [^4] Nvidia A100/RTX3090/V100 and the same initialisation. For task-specific training, we train for 100 epochs with the same learning rate and a batch size set to 16. We choose the final checkpoint based on validation dataset performance (for language adapters, we evaluate on a held-out subset of Wikipedia).
Downstream evaluation is performed on three tasks: Named Entity Recognition (NER), Natural Language Inference (NLI) and Question Answering (QA), covering jointly 13 languages555ar, bg, de, el, es, fr, hi, ru, sw, tr, ur, vi, zh; XQuAD does not cover 4 languages (bg, fr, sw, ur), while the training - to perform cross-lingual evaluation - is performed on the English data. For the NER task, we use the WikiANN [35] dataset and for NLI - XNLI [10]. The QA evaluation is done on XQuAD [5] (we split data 50/50 into valid/test datasets), and the training uses SQuAD 1.1 [36]. Additionally, to evaluate a low-resource scenario for a language not covered during MLLM pre-training, we leverage the Assamese subset from IndicXNLI [1].
4.2 Zero-shot evaluation
Table 1: Zero-shot results of language arithmetic (LA) per language, where one side of the arithmetic equation is the English adapter, and the other is related to the evaluated language adapter (e.g. Italian for Spanish evaluation). As baselines, we report using both adapters in isolation (pEA - proxy via English adapter, pRA - proxy via a related language adapter).
ar bg de el es fr hi ru sw tr ur vi zh AVG
NER (F1 score)
mBERT pEA 33.45 75.62 78.10 69.80 73.42 78.25 63.03 61.63 63.14 72.56 28.72 68.81 36.99 61.81
pRA 33.54 74.51 76.76 69.70 71.53 78.37 58.01 62.05 61.85 68.71 25.08 67.44 33.30 60.07
$LA_{0s}(en,REL)$ 37.22 75.83 77.89 73.42 73.42 78.51 64.11 62.95 64.80 72.56 30.39 69.25 37.65 62.77
[…]
pRA 39.07 75.24 73.12 65.00 67.01 72.74 66.43 58.14 57.53 64.60 36.72 55.82 18.91 57.72
$LA_{0s}(en,REL)$ 43.19 76.57 74.27 70.58 73.48 74.75 68.64 60.37 63.02 69.45 37.22 60.68 23.44 61.20
NLI (Accuracy)
mBERT pEA 61.66 65.43 68.40 62.08 72.22 71.84 58.34 65.87 46.25 56.93 56.49 68.46 66.73 63.13
pRA 53.87 64.49 66.43 61.14 70.84 70.68 54.47 63.21 43.33 54.43 54.51 67.43 66.63 60.88
$LA_{0s}(en,REL)$ 62.06 65.23 68.42 62.44 71.52 72.20 57.96 65.19 46.51 56.95 56.13 69.28 67.33 63.17
[…]
pRA 67.35 76.01 73.67 70.56 69.92 71.46 67.70 74.05 55.89 68.28 62.93 72.36 71.30 69.34
$LA_{0s}(en,REL)$ 68.66 76.69 73.81 73.03 76.97 76.11 68.74 73.65 60.12 70.16 63.87 72.51 73.15 71.35
QA (F1 score)
mBERT pEA 55.93 - 71.93 51.46 72.39 - 50.65 67.46 - 44.99 - 63.51 54.12 59.16
pRA 41.76 - 71.14 53.03 71.67 - 43.22 66.31 - 43.68 - 65.05 54.97 56.76
$LA_{0s}(en,REL)$ 55.93 - 71.32 53.91 72.61 - 50.40 66.36 - 44.64 - 64.16 54.46 59.31
[…]
pRA 62.54 - 69.58 64.98 72.60 - 63.00 69.09 - 63.63 - 70.07 59.15 66.07
$LA_{0s}(en,REL)$ 64.87 - 72.69 71.69 74.45 - 64.01 70.11 - 64.98 - 71.55 61.49 68.43
[…]
Table 1 presents the results of the zero-shot experiment. Language arithmetic consistently outperforms the proxy baselines, reaching up to an improvement of over $3$ F1 points in the XLM-R NER setup and almost $1.5$ F1 for QA. Additionally, we looked at the $\lambda$ parameter derived on a validation dataset. While most cases set the value to over $0.5$ (i.e. preferring the English side, given $LA_{0s}(en,REL)$), the preferred values did not showcase any consistency and pattern. We analyse that phenomenon in Section 5.1.
4.3 Language Arithmetic evaluation
This evaluation assumes that the target language adapter exists in the adapter pool. We test two cases, i.e. $LA(T,en)$ and $LA(T,REL)$, where $REL$ is once again a related language and $T$ is the target language. The results are presented in Table 2. Compared to the baseline direct application of a target language adapter (i.e. MAD-X method), the gains are not as significant as in the case of the zero-shot scenario. Given that a target language adapter is trained on a significant corpus, it gives less room for improvement (this is not the case in the low-resource regimes, as shown in the following Section). However, considering the cost-to-performance ratio and the minimal fatigue that our post-processing method enforces on existing MAD-X pipelines, we can see a constant but minor gain.
Notably, during the evaluation of XLM-R NLI, we observed that the outcomes remain consistent across different methods (e.g. $67.8$ for Swahili). This behaviour is a built-in property of the proposed method. The method has a fail-safe mechanism, which keeps the extreme values during the evaluation of $\lambda\in[0,1]$. Suppose the method does not find a better lambda different than the extreme values (i.e. $\lambda\in\{0,1\}$). In that case, it means that we are using a proxy language directly, and the language arithmetic did not provide an improvement (but kept the final result the same).
Table 2: Results over all tested languages of the improvement of a target language adapter (denoted as $T$, i.e. the MAD-X method [29]) via language arithmetic. We use both English ($en$) and a related language ($REL$) as a right side of a $LA$ function. Additionally, we report a mix ($MIX$) results where we choose better arithmetic combination for each language.
$T\rightarrow$ ar bg de el es fr hi ru sw tr ur vi zh AVG
NER (F1 score)
mBERT $T$ 37.03 79.30 78.50 74.79 74.83 78.32 68.50 65.95 73.65 71.70 40.68 72.64 36.77 65.59
$LA(T,en)$ 37.03 80.11 78.50 75.56 74.64 78.24 66.13 65.95 74.72 72.37 40.68 73.12 39.62 65.90
$LA(T,REL)$ 37.03 79.64 78.50 75.12 74.83 78.26 68.09 66.04 74.50 71.89 40.68 72.88 39.21 65.90
$LA(T,MIX)$ 37.03 80.11 78.50 75.56 74.83 78.26 68.09 66.04 74.72 72.37 40.68 73.12 39.62 66.07
[…]
$LA(T,en)$ 51.32 79.60 77.56 78.65 78.86 76.90 72.27 62.37 74.84 76.28 63.96 66.68 27.55 68.22
$LA(T,REL)$ 51.23 80.16 77.59 78.50 79.06 76.70 71.72 62.34 74.53 76.28 63.96 66.18 28.10 68.18
$LA(T,MIX)$ 51.32 80.16 77.59 78.65 79.06 76.90 72.27 62.37 74.84 76.28 63.96 66.68 28.10 68.32
NLI (Accuracy)
mBERT $T$ 63.41 67.35 67.88 65.23 71.72 72.08 61.24 66.01 58.98 62.99 58.54 68.16 67.64 65.48
$LA(T,en)$ 63.41 67.35 68.46 65.23 72.24 72.24 61.24 66.01 59.16 62.81 58.10 68.54 68.22 65.62
$LA(T,REL)$ 63.41 67.35 68.80 65.09 71.12 72.28 61.08 65.65 58.96 62.99 58.48 68.94 67.90 65.54
$LA(T,MIX)$ 63.41 67.35 68.80 65.23 72.24 72.28 61.24 66.01 59.16 62.99 58.48 68.94 68.22 65.72
[…]
$LA(T,en)$ 70.36 76.59 75.37 74.33 77.41 77.03 69.34 74.73 67.80 71.54 66.03 74.19 73.47 72.94
$LA(T,REL)$ 71.32 76.59 75.35 74.97 77.13 76.97 68.84 74.73 67.80 71.12 66.15 73.85 72.71 72.89
$LA(T,MIX)$ 71.32 76.59 75.37 74.97 77.41 77.03 69.34 74.73 67.80 71.54 66.15 74.19 73.47 73.07
QA (F1 score)
mBERT $T$ 54.46 - 71.17 51.14 73.45 - 48.80 66.92 - 49.27 - 65.85 54.10 59.46
$LA(T,en)$ 55.84 - 71.13 53.41 72.83 - 50.36 67.62 - 48.85 - 63.54 54.15 59.75
$LA(T,REL)$ 55.42 - 70.99 54.71 73.16 - 48.80 67.07 - 49.31 - 64.94 54.94 59.93
$LA(T,MIX)$ 55.84 - 71.13 54.71 73.16 - 50.36 67.62 - 49.31 - 64.94 54.94 60.22
[…]
$LA(T,en)$ 67.61 - 72.71 73.44 75.60 - 67.96 72.02 - 66.57 - 74.07 63.24 70.36
$LA(T,REL)$ 68.00 - 71.62 73.66 75.71 - 67.96 72.03 - 67.01 - 74.12 64.85 70.55
$LA(T,MIX)$ 68.00 - 72.71 73.66 75.71 - 67.96 72.03 - 67.01 - 74.12 64.85 70.67
4.4 Low resource evaluation
Figure 2: NER (left) and NLI (right) evaluation of a set of adapters trained on a Wikipedia subset showcases that language arithmetic $LA(T,en)$ (green, dotted line) provides significant gains when compared against direct usage of the adapter (violet, solid line), especially in a very low-resource regime. The x-axis represents the token budget of each trained language adapter.
Training a language adapter might be troublesome for high-resourced languages due to massive corpora requiring significant computational resources. [^6] Although in our experimental setup, we train each adapter for the same number of steps and choose the best checkpoint based on the validation performance, for low-resourced languages, one could apply an early stopping mechanism in a production-level pipeline. On the other hand, in most languages, we lack data to train a strong language adapter, i.e. language-specific corpora might be either too small or unavailable [15]. We investigate whether language arithmetic can help in such cases. We test our solution in three cases and define three (actual and simulated) evaluation scenarios:
  • Assamese (as) - low resource language, additionally not used in the pre-training of mBERT and XLM-R,
  • Swahili (sw) - low resource language, used in the pre-training,
  • French (fr) - high resource, used in the pre-training. We simulate cases from low to high resources.
We train a series of language adapters with different token budgets for each language, from 10k to 10M (or 1B for French; we limit this particular study to the XLM-R model). Afterwards, we compare the usage of such adapters directly against language arithmetic with three adapters (we use $LA(T,en)$, where $T\in\{as,sw,fr\}$).
Figure 2 presents the results of the evaluation performed on the downstream tasks. The most gain is visible in the most challenging scenario, during the evaluation on the Assamese dataset. In this case, the backbone MLLMs did not encounter the language during the pre-training phase. Although the difference becomes less pronounced in the NER task as we approach the limits of available data, there remains a significant margin for NLI - the difference can be explained by the overlap in the corpora (Wikipedia) between NER and language adapter training tasks, following the findings of Gururangan et al. [13]. For Swahili, where the language is part of the pre-training, the flattening effect begins earlier and affects both tasks. Nevertheless, leveraging language arithmetic still yields improvements.
The simulated case of French showcases that even with a relatively weak language adapter (trained on 10k tokens), the language arithmetic can restore existing knowledge and results in high performance for the language. Moreover, comparing the adapters trained with a different token budget, the results remain similar, without significant fluctuations. We believe that this phenomenon happens because the Multilingual Pre-trained Language Model has seen a much higher amount of French in the pre-training procedure than Swahili (over $35$ times more tokens in XLM-R pre-training; moreover, French is in the top $15$ represented languages). Therefore, even undertrained French adapters have a relatively easy task once they are merged with a robust English adapter. In practical terms, this finding allows us to prototype new languages quicker by estimating the possible end product quality or might serve as an intermediate solution (until the full-corpora adapter is trained).
5 Analysis
5.1 Lambda impact
Our study estimates the $\lambda$ parameter with a small step ($0.05$). Depending on multiple variables that include, among others, model and evaluation dataset sizes or number of languages, running $20$ evaluations might be troublesome (especially when using neural-based metrics, e.g. COMET [38]). Therefore, we analysed the potential impact of choosing a suboptimal lambda with a decrease in evaluation count. The breakdown includes a subset of languages on both tasks (using the XLM-R version). We chose the zero-shot scenario where we performed language arithmetic between English and related language adapters.
In Figure 3, we plot the validation scores with the corresponding baselines, that is, the scores of using directly the adapters. The dotted lines are based on $\lambda=0$ or $\lambda=1$ for clarity, meaning we exclusively use the arithmetic equation’s left or right side (i.e., a specific language). In most cases, a subset of valid $\lambda$ values would improve over the baselines. Moreover, the analysis reveals that a coarser evaluation (e.g., with a step of $0.1$ or $0.2$) would be sufficient, reducing the required number of performed tests up to four times while maintaining most of the improvement. At last, setting the default $\lambda=0.5$ would be near optimum for the analysed subset.
Figure 3: Interpolation of $\lambda$ values for the zero-shot XLM-R scenario (NER - top, NLI - bottom; for QA see Appendix A) on the validation dataset. The horizontal dashed lines represent the baseline scores for both languages used in language arithmetic.
5.2 Language Relatedness
Relatedness of languages is a difficult-to-define concept. At times, in our proposed framework, we might face a choice of multiple, seemingly equally related languages to use for the arithmetic operation. In this analysis, we decided to look at this aspect via a glance at Romance languages. We trained an additional subset of language adapters and formed a pool of 6 languages: Catalan, French, Italian, Portuguese, Romanian and Spanish. Afterwards, we evaluated languages shared in our NER and NLI tasks (Spanish and French) by arithmetic with the entire Romance languages pool.
The results are presented in Table 3 and show that given a related language (in this case, defined as coming from the same language family), there might be very little difference in which one we will use. The relative difference between the best and the worst language reaches only a $0.40$ F1 score in the NER task and the same value in terms of accuracy points in NLI. Nevertheless, the more sophisticated and hand-crafted language choice would improve the results from Table 2.
Table 3: Impact of language relatedness on the language arithmetic. We compare different Romance languages as a right side of $LA$ equation, i.e. $l2$ (both tasks use XLM-R model).
$LA(l_{1}\downarrow,l_{2}\rightarrow)$ ca es fr it pr ro
Eval language NER
es 78.77 - 78.96 79.06 78.91 78.66
fr 76.63 76.92 - 76.70 76.80 76.80
[…]
es 77.07 - 77.23 77.13 77.07 76.83
fr 76.77 76.99 - 76.97 76.95 77.07
5.3 Language vs task vectors
Task vectors exhibit high sparsity and orthogonality, as Ilharco et al. [19] observed. While the former characteristic can be denoted in language vectors (Figure 4), the latter displays different properties, in contrast to task vectors. In Figure 5, we visualise the cosine similarity between evaluated language vectors of language adapters. Notably, the minimal cosine similarity ($0.2$) surpasses the maximum ($0.18$) reported by previous research in the task space [19]. Interestingly, most pairs in the task space oscillate within the range of $0.01$ to $0.03$. At the same time, language vectors surpass $0.2$ in each case, indicating that the orthogonality aspect is an inherent property of task adapters.
Figure 4: Language vectors, similar to task vectors, are extremely sparse. The kernel density estimate plot presents the weights of a Spanish mBERT adapter. The behaviour is consistent across sampled layers and languages.
Based on the cosine similarity observation, we investigated one of the recent task arithmetic extensions, Ties-Merging [44]. This work introduces a three-step algorithm that prevents different parameter interferences, improving upon task arithmetic. The algorithm decreases the cosine similarity via a pruning step and alignment of parameter signs to perform arithmetic only on relevant parameters to the merged tasks. On the experimental details note, as Ties-Merging operates on averaging, not addition, we utilise a different lambda range during validation (as suggested by Yadav et al. [44], $\lambda\in[0.8,1.8]$), and we set Top-Kto the default value of $20$ and additionally to $80$.
Figure 5: Cosine similarity between language vectors of language adapters.
We report the comparison in the zero-shot setting on the NER task (XLM-R version) in Table 4. The Ties-Merging decreases the results significantly; it is not only outperformed by language arithmetic in every language pair but also by other baselines (see Table 1). Moreover, we note that the pruning operation has the reverse effect; higher pruning (i.e. keeping Top-Klower) decreases the performance (in contrast to task vectors) by making language vectors more sparse and, hence, closer to orthogonal.
Table 4: Ties-Merging evaluation in the zero-shot setup on the NER task (XLM-R version). In the case of language arithmetic, where the language vectors have a higher overlap (i.e. higher cosine similarity), removing parameter interference decreases the overall performance.
Method AVG F1 score
LA 61.20
Ties-Merging (Top-K20) 56.40
Ties-Merging (Top-K80) 59.88
One interpretation of the phenomena can be the different goals of the arithmetic: in the multi-task setup, we try to include multiple, often disconnected, tasks into a single task vector. In contrast, the language vectors’ goal is to include the knowledge of the closely related language rather than remove the harmful artefacts. Our findings indicate that language arithmetic has different characteristics than task arithmetic, and the follow-up works that improve upon task arithmetic might not be suited for the multilingual context.
6 Related Work
Knowledge composition from multiple, independently trained adapters has been widely discussed in the literature. However, unlike our work, the solutions require substantial changes to the vanilla adapter setup. The previous work either requires additional parameters to learn a parameterised composition function/a gating module to combine/steer the flow through the suitable adapter(s), or needs a specific training procedure that increases the complexity of the overall solution or, in most cases, both [33, 30, 22, 26, 9, 20, 41]. Moreover, to prevent specifically negative interference, hyper-adapters [7] were proposed using hyper-networks [14], and Ansell et al. [3] applied sparse fine-tuning to compose task and language masks. Unlike the prior studies mentioned earlier, our attempt is training-free and does not modify the base architecture. The most conceptually similar work is proposed by Chronopoulou et al. [8]; however, they operate on the notion of sample similarity to a subset of domains in a domain adaptation regime. Additionally, concurrent to our work, Parović et al. [27] show intial potential of task arithemtic in cross-lingual transfer based on a full fine-tuning setup. At last, we denote the rise of task arithmetic use cases, e.g. vision tasks or cross-task generalisation [39, 18].
7 Conclusion
We have proposed language arithmetic, which enhances language adapters based on task arithmetic learning via addition. It is a training-free method and functions as a post-processing technique for the adapters. Our experiments have shown that LA is particularly beneficial in zero-shot and low-resourced cases. At last, we highlight the differences between language and task arithmetic.
In our future work, we plan to extend language arithmetic by incorporating additional components into the sum. Additionally, we aim to adapt other elements of the task arithmetic framework, initially designed for tasks, i.e. task analogies and forgetting via negation, to a multilingual setup with an analysis of the differences between multi-task and multilingual arithmetic context. Furthermore, we will evaluate LA’s performance on various non-classification tasks.
8 Limitations
Our work was tested on English-centric task training and could be extended to different languages with more PEFT methods. Moreover, applying multi-source training based on the work of Ansell et al. [4] could provide better robustness of the task adapters and a more thorough analysis. At last, our analysis of the differences between task and language vectors is preliminary, without definitive conclusions.
Appendix A Lambda impact - QA
Figure 6 presents the analysis of lambda impact for the Question Answering task. For details, refer to Section 5.1.
Figure 6: Interpolation of $\lambda$ values for the zero-shot QA XLM-R scenario on the validation dataset. The horizontal dashed lines represent the baseline scores for both languages used in language arithmetic.
Appendix B Related languages choice
We derived related languages for the language adapters based on the available pool, i.e. we reused the adapters from other languages from the task (we trained two bonus ones, Persian and Italian, apart from the analysis in Section 5.2). While the related language could have been more suitable in a few cases, there is always a trade-off between relatedness, criterion for relatedness and available corpora. Moreover, the perfectly related language matching would be an unfair comparison as it would assume the availability of the ideal adapter’s pool, which is unlikely in practice. Table 5 presents the related language choice used in our experiments.
Table 5: Languages used in the experiments with corresponding related languages. In the most cases, we used a related language from the already existing pool (i.e. evaluated languages, the first rows labelled “Lang.”); however, we trained some additional ones, including Italian and Persian.
Lang. ar bg de el es fr hi
Related Lang. fa ru fr de it it ur
[…]
Related Lang. bg ar de hi zh vi

### v2

@@
 In response, we introduce a novel method called language arithmetic, which enables training-free post-processing to address this limitation.
-Inspired by the task arithmetic framework, we apply learning via addition to the language adapters, transitioning the framework from a multi-task to a multilingual setup.
+Extending the task arithmetic framework, we apply learning via addition to the language adapters, transitioning the framework from a multi-task to a multilingual setup.
 The effectiveness of the proposed solution is demonstrated on three downstream tasks in a MAD-X-based set of cross-lingual schemes, acting as a post-processing procedure.
-Language arithmetic consistently improves the baselines with significant gains in the most challenging cases of zero-shot and low-resource applications.
+Language arithmetic consistently improves the baselines with significant gains, especially in the most challenging case of zero-shot application.
 Our code and models are available at https://github.com/mklimasz/language-arithmetic.
+No Train but Gain:
+Language Arithmetic for training-free Language Adapters enhancement
+Mateusz Klimaszewski1,2, Piotr Andruszkiewicz1, Alexandra Birch2 1Warsaw University of Technology, 2University of Edinburgh Correspondence: mklimasz@ed.ac.uk
 1 Introduction
-The recent progress of large language models has led to a huge interest in how well they can perform not just in English but across multiple languages and invoked the rise of Multilingual Pre-trained Language Models (MLLMs) [11, 43, 2].
+The recent progress of large language models has raised the question of how well they perform not just in English but across multiple languages which has spurred interest in Multilingual Pre-trained Language Models (MLLMs) Conneau et al.
+(2020); Workshop (2023); Alves et al.
+(2024).
 These models serve as general-purpose solutions that can be adapted and applied to various Natural Language Processing tasks.
-Notably, MLLMs demonstrate zero-shot cross-lingual capabilities, allowing them to generalise effectively to downstream tasks even when trained in a language different from the target language.
-The positive transfer of abilities from both related languages and high-quality training data from unrelated languages has meant that MLLMs have reported state-of-the-art performance in low-resourced languages [24].
-However, this benefit does not always extend to high-resourced languages [21].
-In such cases, the quality of MLLMs tends to decrease compared to their monolingual counterparts [25, 23, among others] due to negative interference phenomena [42].
-Additionally, the curse of multilinguality [11] reveals the existence of a trade-off between language coverage and model capacity.
+Notably, MLLMs demonstrate zero-shot cross-lingual capabilities, allowing them to generalise effectively to downstream tasks even when pre-trained in a language different from the target language.
+The positive transfer of abilities from both related languages and high-quality training data from unrelated languages has meant that MLLMs have reported state-of-the-art performance in low-resourced languages Muennighoff et al.
+(2023).
+However, this benefit does not always extend to high-resourced languages Kocmi et al.
+(2021).
+In such cases, the quality of MLLMs tends to decrease compared to their monolingual counterparts (Nozza et al., 2020; Martin et al., 2020, among others) due to negative interference phenomena Wang et al.
+(2020).
+Additionally, the curse of multilinguality Conneau et al.
+(2020) reveals the existence of a trade-off between language coverage and model capacity.
 Consequently, MLLMs must carefully limit the number of languages included during the pre-training phase.
-Modular deep learning (MDL) [32] methods come to the rescue to avoid the abovementioned limitations, enabling the extension of MLLMs to support any number of languages [6, 40, 34, 29, 31].
+Modular deep learning (MDL) Pfeiffer et al.
+(2023) methods help to avoid negative interference and limited model capacity , enabling the extension of MLLMs to support any number of languages Bapna and Firat (2019); Üstün et al.
+(2020); Philip et al.
+(2020); Pfeiffer et al.
+(2020b, 2022).
 MDL methods adapt the model to arbitrary tasks and languages by isolating components from each other (and the backbone MLLM) via parameter-efficient extensions.
-Examples of parameter-efficient modules are LoRA [17] and adapters [37, 16], which are low-budget (in terms of parameters) bottleneck layers that increase an MLLM size by just a fraction.
-Language adapters especially allow the modularisation of language-specific knowledge by training on a raw, unlabelled corpus in a target language.
+Examples of parameter-efficient modules are adapters Rebuffi et al.
+(2017); Houlsby et al.
+(2019), which are low-budget (in terms of parameters) bottleneck layers that increase an MLLM size by just a fraction.
+Language adapters Pfeiffer et al.
+(2020b) allow the modularisation of language-specific knowledge by training on a raw, unlabelled corpus for specific languages.
 The limitation of the MDL and language adapters is their isolation.
 While they lift the curse of multinguality and prevent negative interference, at the same time, language adapters limit the possible impact of positive transfer.
-Previous attempts to address these challenges — such as training bilingual [26] or language-family [9] adapters — do not scale effectively.
+Previous attempts to address these challenges — such as training bilingual Parović et al.
+(2022) or language-family Chronopoulou et al.
+(2023b) adapters — do not scale effectively.
 In our work, we tackle this limitation as a post-processing step.
-Leveraging recent insights from task arithmetic [19], specifically learning via addition, we augment language adapters with missing related language knowledge – a concept we term language arithmetic.
-Remarkably, this training-free approach enhances language adapters across various scenarios, from zero-shot settings to low-resource regimes.
+Leveraging recent insights from task arithmetic Ilharco et al.
+(2023), specifically learning via addition, we augment language adapters with missing related language knowledge – a concept we term language arithmetic.
+Remarkably, this training-free approach can enhance not only existing language adapters but also offer zero-shot performance.
 To summarise, our contributions are as follows:
   • A novel training-free post-processing method named language arithmetic that enhances language adapters.
-  • We conduct a cross-lingual evaluation on three downstream tasks and Multilingual Pre-trained Language Models with test cases that include zero-shot and low-resource setups.
-  • We provide an analysis of language arithmetic internal components.
+  • We conduct a cross-lingual evaluation on three downstream tasks (NER, NLI and QA) and two Multilingual Pre-trained Language Models (XLM-R, mBERT) with test cases that include zero-shot and low-resource setups in a diverse group of 13 languages.
+  • We provide an analysis of language arithmetic internal components (including a comparison with task arithmetic) and show improvement up to 3 F1 points without any additional training involved.
 2 Background
-Our research builds upon the contributions of Ilharco et al.
-[19] and Zhang et al.
-[45].
+Our research builds upon the task arithmetic contributions of Ilharco et al.
+(2023) and Zhang et al.
+(2023).
 The following Section provides the background and serves as a gentle introduction to the concept of task vectors and task arithmetic.
@@
 Moreover, vectors from different tasks are typically close to orthogonal, and Ilharco et al.
-[19] speculate that this enables the combination of task vectors via addition with minimal interference.
-In the Parameter-Efficient Fine-tuning (PEFT) regime, we can reduce a task vector to only newly introduced PEFT parameters [45].
-Another way of looking at this is a task vector that is a zero-valued vector for all the parameters apart from PEFT weights, i.e.
-a task vector would not modify the base, pre-trained checkpoint.
+(2023) speculate that this enables the combination of task vectors via addition with minimal interference.
+In our work, we focus on parameter-efficient fine-tuning (PEFT) and use language adapters.
+Therefore, we reduce the task vector and underlying model weights represented by $\theta$ to newly introduced parameters (i.e.
+we exclude the backbone MLLM, which is frozen across all the models, following the work of Zhang et al.
+(2023))
 3 Method
 We propose language arithmetic that transitions the task arithmetic concept from a multi-task to a multilingual setup.
-In this Section, we describe the language arithmetic alongside its application as a training-free, post-processing step to a MAD-X cross-lingual framework [29].
+In this Section, we describe the language arithmetic alongside its application as a training-free, post-processing step to a MAD-X cross-lingual framework Pfeiffer et al.
+(2020b).
 3.1 Language arithmetic
@@
 This approach means that instead of merging downstream tasks, we target a problem of cross-lingual performance.
-We propose to apply learning via addition to languages, and in Section 5.3, we demonstrate the discrepancies when comparing language and task vectors.
-Our study focuses specifically on the language adapters [34, 29].
+We propose to apply learning via addition to languages, and in Appendix E, we demonstrate the discrepancies when comparing language and task vectors.
+Our study focuses specifically on the language adapters Philip et al.
+(2020); Pfeiffer et al.
+(2020b).
 Due to overlapping abbreviations, we use the LA exclusively as the former, i.e.
@@
 Throughout the paper, the equation above is abbreviated as a function: $LA(en,es)$ with lambda as a default parameter.
-Whenever we use language arithmetic in a zero-shot context, we add a subscript to the function name, i.e.
-$LA_{0s}(en,es)$ means that language arithmetic between English and Spanish is applied to a different language, e.g.
-French.
+Additionally, for clarity reasons, we denote target language in a subscript to distinguish different use-cases.
+For example, $LA_{fr}(en,es)$ means that language arithmetic between English and Spanish is applied to a different language - in this case French (zero-shot application), or $LA_{es}(en,es)$ meaning that the target is Spanish (non-zero-shot use-case, due to Spanish being a part of the LA equation).
+Figure 1: Language arithmetic as an extension of the MAD-X framework.
+Given language and task adapters (left), language arithmetic (right) enables post-processing, training-free improvement in two use-cases: (i) zero-shot where a language adapter for a target language was not trained (presented in the figure as Spanish, which was not part of existing language adapters pool, $LA_{es}(en,fr)$) or (ii) to improve existing language adapters via arithmetic with either related language or a language on which task adapter was trained (e.g.
+$LA_{fr}(en,fr)$).
 Language arithmetic is a training-free method, taking advantage of already pre-trained modules.
@@
 3.2 Application
-We apply the language arithmetic as an enhancement method for language adapters.
-We assume two scenarios: (i) zero-shot case, where we do not have a language adapter in the desired language and (ii) “regular” enhancement case, where we improve the existing language adapter (in high- and low-resourced languages).
-We evaluate our post-processing method as an extension of the MAD-X framework [29] to challenge our solution in a cross-lingual manner. The overview of the schema is presented in Figure 1.
-[…]
-The MAD-X consists of the following:
+We evaluate our post-processing method as an extension of the MAD-X framework Pfeiffer et al.
+(2020b) to challenge our solution in a cross-lingual manner.
+The overview of the schema is presented in Figure 1.
+The MAD-X consists of the following steps:
   1. 1.
-    Choosing MLLM
+    Training language adapter(s)
   2. 2.
-[…]
+    Training task adapter
   3. 3.
-[…]
+    Cross-lingual inference
+In our work, we introduce an additional step:
   1. 4.
-[…]
-In our work, we introduce an additional step:
-  1. 5.
-[…]
-3.2.1 MLLM
-The first step usually assumes the availability of a Multilingual Pre-trained Language Model. In our work, we focus on mBERT111bert-base-multilingual-cased [12] and XLM-R222xlm-roberta-base [11].
-3.2.2 Training language & task adapters
-The following two steps depend on the downstream dataset.
-In the most extreme case, we might have a dataset in one specific language and need to take advantage of cross-lingual performance.
-Therefore, step number 2 trains a language adapter on raw corpora corresponding to the task language using regular masked language modelling loss in a self-supervised manner.
-During this step, the MLLM is frozen.
-The third step freezes both the backbone and language adapter and trains a task adapter on a downstream task dataset.
-Given a set of tasks or if a new task appears, we can repeat this step as long as a required language adapter exists, i.e.
+    Post-processing via language arithmetic
+In the following Sections 3.2.1-3.2.4, we present the framework and our proposed post-processing via language arithmetic extension, exploring two use cases: (i) a zero-shot case, where a target language adapter does not exist and (ii) an enhancement case, where we prove existing language adapters (in high- and low-resourced languages).
+3.2.1 Training language adapter(s)
+In the first step, the MAD-X framework trains language adapters.
+These adapters are trained on raw corpora using masked language modelling loss in a self-supervised manner.
+The MLLM is frozen during this step, and we only optimise the newly introduced adapter.
+The training must be done for languages corresponding to the downstream tasks (e.g.
+if we have an English NER dataset, we need an English language adapter, apart from other desired target languages).
+Additionally, the adapters form a pool that is leveraged during cross-lingual inference.
+3.2.2 Training task adapters
+The following step freezes a backbone MLLM and a language adapter and trains a task adapter on a downstream task dataset.
+Given a set of tasks or if a new task appears, we can repeat this step as long as the required language adapter exists in the available pool, i.e.
 a language adapter that matches the task’s language.
-3.2.3 Additional language adapters
-The last, fourth step is the heart of the MAD-X framework.
-We can now train language adapters in any language, leverage the pre-trained task adapter(s) and obtain a cross-lingual performance by connecting a new language adapter with a task adapter (i.e.
-routing first via language adapter and then task) to obtain language-task pair unseen during training.
-The growing pool of pre-trained adapters can be accessed at public repositories like AdapterHub [28] and reused for further use cases.
+3.2.3 Cross-lingual inference
+Having trained a task adapter, we can leverage a pool of pre-trained language adapters and obtain a cross-lingual performance by connecting any existing language adapter with a newly trained task adapter (i.e.
+routing first via language adapter and then task adapter).
+The growing pool of pre-trained adapters can be accessed at public repositories like AdapterHub Pfeiffer et al.
+(2020a) and reused for further use cases.
 3.2.4 Post-processing via language arithmetic
-First, we assume the use-case where the pool of language adapters does not contain the target languages, i.e.
-zero-shot scenario.
-In this study, we add English as a task-specific language to a language related to the target.
-In the second language adapter enhancement scenario, we apply language arithmetic in its default form, where the goal is to enhance the existing target language adapter (via either English or a related language).
-At last, we investigate (simulated and actual) low-resource cases.
+Our method builds upon the MAD-X framework in two enhancement scenarios.
+First, we assume a situation where the pool of language adapters does not contain a desired target language, i.e.
+a zero-shot scenario.
+In contrast to the previous works that try to improve existing adapters, this use-case is an alternative to routing via either a related language or a task language (here, by task language, we understand the language on which the task adapter was trained, in contrast to a target language - on which we want to evaluate).
+Instead of choosing the better-performing proxy, language arithmetic proposes to combine these two (with better results, as shown in Section 4.2).
+In the second language adapter enhancement scenario, we apply language arithmetic as a more common goal, trying to improve existing language adapters; however, our method does that without any training.
+Here, we combine the existing target language adapter with either a related language (we define related languages in Section 4.1.1) or, once again, a task language.
 4 Experiments
 4.1 Experimental setup
-We use the AdapterHub library [28] for all our experiments.
-For language adapters, we train on the Wikipedia corpora33320331101.xx checkpoint https://huggingface.co/datasets/wikimedia/wikipedia for 250k steps with a learning rate of 1e-4, an effective batch size of 64 using a single GPU [^4] Nvidia A100/RTX3090/V100 and the same initialisation.
+4.1.1 Datasets
+Downstream evaluation is performed on three tasks: Named Entity Recognition (NER), Natural Language Inference (NLI) and Question Answering (QA), covering jointly 13 languages111ar, bg, de, el, es, fr, hi, ru, sw, tr, ur, vi, zh; XQuAD does not cover 4 languages (bg, fr, sw, ur), while the training - to perform cross-lingual evaluation - is performed on the English data.
+For the NER task, we use the WikiANN Rahimi et al.
+(2019) dataset and for NLI - XNLI Conneau et al.
+(2018).
+The QA evaluation is done on XQuAD Artetxe et al.
+(2020) (we split data 50/50 into valid/test datasets), and the training uses SQuAD 1.1 Rajpurkar et al.
+(2016).
+Additionally, to evaluate a low-resource scenario for a language not covered during MLLM pre-training, we leverage the Assamese subset from IndicXNLI Aggarwal et al.
+(2022).
+Related languages
+To automatically establish a related language needed for language arithmetic, we query URIEL and lang2vec library Littell et al.
+(2017); Malaviya et al.
+(2017).
+During the related language query, we limited the options to 13 downstream task languages for which we had already pre-computed language adapters.
+This limitation means that for some languages, we would be able to find a stronger performing pairing and that the ceiling for our method is higher than the presented (we denote that the performance of hypothetical paring would also depend on data availability, i.e.
+a paired language must be not only related but also have representative corpora; based on this we show analysis and improved performance in Section 5.2).
+However, considering the limitations of the lang2vec, we decided to keep this simplification.
+At last, a language can have a set of equally good related languages.
+Therefore, in practical terms, it is not feasible for our study to train all possible options for each language - our simplification stands as a reasonable, real-world proxy.
+We provide the list of related languages in the Appendix A.
+4.1.2 Implementation & training
+In our work, we focus on two of the most popular multilingual PLMs [^2] According to downloads from Huggingface hf.co/models?language=multilingual&sort=downloads: mBERT333bert-base-multilingual-cased Devlin et al.
+(2019) and XLM-R444xlm-roberta-base Conneau et al.
+(2020).
+We implement our method using the AdapterHub library Pfeiffer et al.
+(2020a).
+For language adapters, we train on the Wikipedia corpora55520331101.xx checkpoint hf.co/datasets/wikimedia/wikipedia for 250k steps with a learning rate of 1e-4, an effective batch size of 64 using a single GPUand the same initialisation.
 For task-specific training, we train for 100 epochs with the same learning rate and a batch size set to 16.
 We choose the final checkpoint based on validation dataset performance (for language adapters, we evaluate on a held-out subset of Wikipedia).
-Downstream evaluation is performed on three tasks: Named Entity Recognition (NER), Natural Language Inference (NLI) and Question Answering (QA), covering jointly 13 languages555ar, bg, de, el, es, fr, hi, ru, sw, tr, ur, vi, zh; XQuAD does not cover 4 languages (bg, fr, sw, ur), while the training - to perform cross-lingual evaluation - is performed on the English data.
-For the NER task, we use the WikiANN [35] dataset and for NLI - XNLI [10].
-The QA evaluation is done on XQuAD [5] (we split data 50/50 into valid/test datasets), and the training uses SQuAD 1.1 [36].
-Additionally, to evaluate a low-resource scenario for a language not covered during MLLM pre-training, we leverage the Assamese subset from IndicXNLI [1].
+In our main experiments, we report the scores as an average over three independent runs with different seeds (for both language and task adapters).
+Additionally, to improve efficiency and reduce GPU memory utilisation, we adopt a half-precision (FP16) setting.
 4.2 Zero-shot evaluation
-Table 1: Zero-shot results of language arithmetic (LA) per language, where one side of the arithmetic equation is the English adapter, and the other is related to the evaluated language adapter (e.g.
-Italian for Spanish evaluation).
-As baselines, we report using both adapters in isolation (pEA - proxy via English adapter, pRA - proxy via a related language adapter).
-ar bg de el es fr hi ru sw tr ur vi zh AVG
-NER (F1 score)
-mBERT pEA 33.45 75.62 78.10 69.80 73.42 78.25 63.03 61.63 63.14 72.56 28.72 68.81 36.99 61.81
-pRA 33.54 74.51 76.76 69.70 71.53 78.37 58.01 62.05 61.85 68.71 25.08 67.44 33.30 60.07
-$LA_{0s}(en,REL)$ 37.22 75.83 77.89 73.42 73.42 78.51 64.11 62.95 64.80 72.56 30.39 69.25 37.65 62.77
-[…]
-pRA 39.07 75.24 73.12 65.00 67.01 72.74 66.43 58.14 57.53 64.60 36.72 55.82 18.91 57.72
-$LA_{0s}(en,REL)$ 43.19 76.57 74.27 70.58 73.48 74.75 68.64 60.37 63.02 69.45 37.22 60.68 23.44 61.20
-NLI (Accuracy)
-mBERT pEA 61.66 65.43 68.40 62.08 72.22 71.84 58.34 65.87 46.25 56.93 56.49 68.46 66.73 63.13
-pRA 53.87 64.49 66.43 61.14 70.84 70.68 54.47 63.21 43.33 54.43 54.51 67.43 66.63 60.88
-$LA_{0s}(en,REL)$ 62.06 65.23 68.42 62.44 71.52 72.20 57.96 65.19 46.51 56.95 56.13 69.28 67.33 63.17
-[…]
-pRA 67.35 76.01 73.67 70.56 69.92 71.46 67.70 74.05 55.89 68.28 62.93 72.36 71.30 69.34
-$LA_{0s}(en,REL)$ 68.66 76.69 73.81 73.03 76.97 76.11 68.74 73.65 60.12 70.16 63.87 72.51 73.15 71.35
-QA (F1 score)
-mBERT pEA 55.93 - 71.93 51.46 72.39 - 50.65 67.46 - 44.99 - 63.51 54.12 59.16
-pRA 41.76 - 71.14 53.03 71.67 - 43.22 66.31 - 43.68 - 65.05 54.97 56.76
-$LA_{0s}(en,REL)$ 55.93 - 71.32 53.91 72.61 - 50.40 66.36 - 44.64 - 64.16 54.46 59.31
-[…]
-pRA 62.54 - 69.58 64.98 72.60 - 63.00 69.09 - 63.63 - 70.07 59.15 66.07
-$LA_{0s}(en,REL)$ 64.87 - 72.69 71.69 74.45 - 64.01 70.11 - 64.98 - 71.55 61.49 68.43
-[…]
-Table 1 presents the results of the zero-shot experiment.
-Language arithmetic consistently outperforms the proxy baselines, reaching up to an improvement of over $3$ F1 points in the XLM-R NER setup and almost $1.5$ F1 for QA.
-Additionally, we looked at the $\lambda$ parameter derived on a validation dataset.
+The zero-shot evaluation assumes a scenario where the language adapter pool does not contain a desired target language (e.g.
+lack of Spanish in Figure 1).
+The baselines are based on routing, i.e.
+we proxy either by an English adapter (proxy via a task language, as the task adapter was trained using English data) or a related language (e.g.
+French for Spanish).
+Language arithmetic serves a solution that, instead of choosing a better proxy, combines the adapter’s tuple: $LA_{t}(en,rel)$, where $rel$ symbolises a related language and $t$ stands for a target language (e.g.
+$LA_{es}(en,fr)$).
+Figure 2: Zero-shot XLM-R language arithmetic evaluation, where one side of the arithmetic is an English adapter, and the other is related to the target language adapter (e.g.
+French for Spanish - $LA_{es}(en,fr)$).
+The values above bars present a relative difference to a better proxy.
+See Figure 6 for the mBERT model.
+Figures 2 and 6 present the results of the zero-shot experiment.
+Language arithmetic consistently outperforms the proxy baselines for all the setups, reaching over $3.1$ F1 points improvement in the NER task and $1.1$ F1 for QA (XLM-R).
+These results indicate that language arithmetic is a feasible, low-cost method that one can apply in the lack of an existing target language adapter.
+Additionally, we investigate how the $\lambda$ parameter impacts the downstream evaluation.
+The goal was to understand how much weight is given to English vs related language.
+We looked at the validation performance over different $\lambda$ thresholds.
 While most cases set the value to over $0.5$ (i.e.
-preferring the English side, given $LA_{0s}(en,REL)$), the preferred values did not showcase any consistency and pattern.
-We analyse that phenomenon in Section 5.1.
-4.3 Language Arithmetic evaluation
-This evaluation assumes that the target language adapter exists in the adapter pool.
+preferring the English side, given $LA_{t}(en,rel)$), the preferred values did not showcase any consistency and pattern.
+We analyse this further in Section 5.1.
+4.3 Improving existing language adapters
+This evaluation assumes that a target language adapter exists in the adapter pool.
 We test two cases, i.e.
-$LA(T,en)$ and $LA(T,REL)$, where $REL$ is once again a related language and $T$ is the target language.
-The results are presented in Table 2.
-Compared to the baseline direct application of a target language adapter (i.e.
-MAD-X method), the gains are not as significant as in the case of the zero-shot scenario.
+$LA_{t}(en,t)$ and $LA_{t}(rel,t)$, where $rel$ is once again a related language and $t$ is the target language.
+Additionally, we provide a combination of these two approaches (referred to as $LA_{t}(en/rel,t)$), where for each language, we choose a better pairing (so either $en$ or $rel$).
+This solution resembles a practical compromise between cost and performance and serves as a proxy for the ceiling of our method (discussed in Section 4.1.1).
+The results are presented in Figures 3 and 7.
+Compared to the baseline direct application of a target language adapter (i.e., the MAD-X method), the gains are not as significant as in the case of the zero-shot scenario.
+Moreover, in contrast to the previous Section’s study, MLLMs showcase a different behaviour, as language arithmetic provides less benefit for XLM-R than for the mBERT model (e.g.
+$+0.38$ XLM-R vs $+2.41$ mBERT in F1, QA).
+The drop in performance of language arithmetic compared to the zero-shot use case is not surprising.
 Given that a target language adapter is trained on a significant corpus, it gives less room for improvement (this is not the case in the low-resource regimes, as shown in the following Section).
-However, considering the cost-to-performance ratio and the minimal fatigue that our post-processing method enforces on existing MAD-X pipelines, we can see a constant but minor gain.
-Notably, during the evaluation of XLM-R NLI, we observed that the outcomes remain consistent across different methods (e.g.
-$67.8$ for Swahili).
-This behaviour is a built-in property of the proposed method.
-The method has a fail-safe mechanism, which keeps the extreme values during the evaluation of $\lambda\in[0,1]$.
-Suppose the method does not find a better lambda different than the extreme values (i.e.
-$\lambda\in\{0,1\}$).
-In that case, it means that we are using a proxy language directly, and the language arithmetic did not provide an improvement (but kept the final result the same).
-Table 2: Results over all tested languages of the improvement of a target language adapter (denoted as $T$, i.e.
-the MAD-X method [29]) via language arithmetic.
-We use both English ($en$) and a related language ($REL$) as a right side of a $LA$ function.
-Additionally, we report a mix ($MIX$) results where we choose better arithmetic combination for each language.
-$T\rightarrow$ ar bg de el es fr hi ru sw tr ur vi zh AVG
-NER (F1 score)
-mBERT $T$ 37.03 79.30 78.50 74.79 74.83 78.32 68.50 65.95 73.65 71.70 40.68 72.64 36.77 65.59
-$LA(T,en)$ 37.03 80.11 78.50 75.56 74.64 78.24 66.13 65.95 74.72 72.37 40.68 73.12 39.62 65.90
-$LA(T,REL)$ 37.03 79.64 78.50 75.12 74.83 78.26 68.09 66.04 74.50 71.89 40.68 72.88 39.21 65.90
-$LA(T,MIX)$ 37.03 80.11 78.50 75.56 74.83 78.26 68.09 66.04 74.72 72.37 40.68 73.12 39.62 66.07
-[…]
-$LA(T,en)$ 51.32 79.60 77.56 78.65 78.86 76.90 72.27 62.37 74.84 76.28 63.96 66.68 27.55 68.22
-$LA(T,REL)$ 51.23 80.16 77.59 78.50 79.06 76.70 71.72 62.34 74.53 76.28 63.96 66.18 28.10 68.18
-$LA(T,MIX)$ 51.32 80.16 77.59 78.65 79.06 76.90 72.27 62.37 74.84 76.28 63.96 66.68 28.10 68.32
-NLI (Accuracy)
-mBERT $T$ 63.41 67.35 67.88 65.23 71.72 72.08 61.24 66.01 58.98 62.99 58.54 68.16 67.64 65.48
-$LA(T,en)$ 63.41 67.35 68.46 65.23 72.24 72.24 61.24 66.01 59.16 62.81 58.10 68.54 68.22 65.62
-$LA(T,REL)$ 63.41 67.35 68.80 65.09 71.12 72.28 61.08 65.65 58.96 62.99 58.48 68.94 67.90 65.54
-$LA(T,MIX)$ 63.41 67.35 68.80 65.23 72.24 72.28 61.24 66.01 59.16 62.99 58.48 68.94 68.22 65.72
-[…]
-$LA(T,en)$ 70.36 76.59 75.37 74.33 77.41 77.03 69.34 74.73 67.80 71.54 66.03 74.19 73.47 72.94
-$LA(T,REL)$ 71.32 76.59 75.35 74.97 77.13 76.97 68.84 74.73 67.80 71.12 66.15 73.85 72.71 72.89
-$LA(T,MIX)$ 71.32 76.59 75.37 74.97 77.41 77.03 69.34 74.73 67.80 71.54 66.15 74.19 73.47 73.07
-QA (F1 score)
-mBERT $T$ 54.46 - 71.17 51.14 73.45 - 48.80 66.92 - 49.27 - 65.85 54.10 59.46
-$LA(T,en)$ 55.84 - 71.13 53.41 72.83 - 50.36 67.62 - 48.85 - 63.54 54.15 59.75
-$LA(T,REL)$ 55.42 - 70.99 54.71 73.16 - 48.80 67.07 - 49.31 - 64.94 54.94 59.93
-$LA(T,MIX)$ 55.84 - 71.13 54.71 73.16 - 50.36 67.62 - 49.31 - 64.94 54.94 60.22
-[…]
-$LA(T,en)$ 67.61 - 72.71 73.44 75.60 - 67.96 72.02 - 66.57 - 74.07 63.24 70.36
-$LA(T,REL)$ 68.00 - 71.62 73.66 75.71 - 67.96 72.03 - 67.01 - 74.12 64.85 70.55
-$LA(T,MIX)$ 68.00 - 72.71 73.66 75.71 - 67.96 72.03 - 67.01 - 74.12 64.85 70.67
+This is also a potential explanation of a different behaviour between the evaluated MLLMs, considering the overall more robust performance of XLM-R over mBERT.
+However, considering the cost-to-performance ratio and the minimal fatigue that our post-processing method enforces on existing MAD-X pipelines, we can see a constant gain on average across all the experiments and training runs.
+Figure 3: Variants of language arithmetic compared to the MAD-X method in the use-case to improve an existing target language adapter.
+The values above bars present a difference between a better LA setup and the MAD-X framework for the XLM-R model (see Figure 7 for mBERT).
 4.4 Low resource evaluation
-Figure 2: NER (left) and NLI (right) evaluation of a set of adapters trained on a Wikipedia subset showcases that language arithmetic $LA(T,en)$ (green, dotted line) provides significant gains when compared against direct usage of the adapter (violet, solid line), especially in a very low-resource regime.
+((a)) NER
+((b)) NLI
+Figure 4: NER and NLI evaluation of a set of adapters trained on a Wikipedia subset showcases that language arithmetic $LA_{t}(t,en)$ (green, dotted line) provides significant gains when compared against direct usage of the adapter (violet, solid line), especially in a very low-resource regime.
 The x-axis represents the token budget of each trained language adapter.
@@
 On the other hand, in most languages, we lack data to train a strong language adapter, i.e.
-language-specific corpora might be either too small or unavailable [15].
-We investigate whether language arithmetic can help in such cases.
+language-specific corpora might be either too small or unavailable Haddow et al.
+(2022).
+We investigate whether LA can help in such cases.
 We test our solution in three cases and define three (actual and simulated) evaluation scenarios:
-  • Assamese (as) - low resource language, additionally not used in the pre-training of mBERT and XLM-R,
+  • Assamese (as) - low resource language, additionally not used in the pre-training of a base MLLM,
   • Swahili (sw) - low resource language, used in the pre-training,
@@
 We train a series of language adapters with different token budgets for each language, from 10k to 10M (or 1B for French; we limit this particular study to the XLM-R model).
-Afterwards, we compare the usage of such adapters directly against language arithmetic with three adapters (we use $LA(T,en)$, where $T\in\{as,sw,fr\}$).
+Afterwards, we compare the usage of such adapters directly against language arithmetic with three adapters (we use $LA_{t}(t,en)$, where $t\in\{as,sw,fr\}$).
 Figure 4 presents the results of the evaluation performed on the downstream tasks.
@@
 Although the difference becomes less pronounced in the NER task as we approach the limits of available data, there remains a significant margin for NLI - the difference can be explained by the overlap in the corpora (Wikipedia) between NER and language adapter training tasks, following the findings of Gururangan et al.
-[13].
+(2020).
 For Swahili, where the language is part of the pre-training, the flattening effect begins earlier and affects both tasks.
@@
 Moreover, comparing the adapters trained with a different token budget, the results remain similar, without significant fluctuations.
-We believe that this phenomenon happens because the Multilingual Pre-trained Language Model has seen a much higher amount of French in the pre-training procedure than Swahili (over $35$ times more tokens in XLM-R pre-training; moreover, French is in the top $15$ represented languages).
+We believe that this phenomenon happens because the MLLM has seen a much higher amount of French in the pre-training procedure than Swahili (over $35$ times more tokens in XLM-R pre-training; moreover, French is in the top $15$ represented languages).
 Therefore, even undertrained French adapters have a relatively easy task once they are merged with a robust English adapter.
@@
 Our study estimates the $\lambda$ parameter with a small step ($0.05$).
-Depending on multiple variables that include, among others, model and evaluation dataset sizes or number of languages, running $20$ evaluations might be troublesome (especially when using neural-based metrics, e.g.
-COMET [38]).
+This analysis investigates how sensitive this parameter is in the language arithmetic.
+Depending on multiple variables that include model and evaluation dataset sizes or a number of languages, running $20$ evaluations might be costly (especially when using neural-based metrics, e.g.
+COMET Rei et al.
+(2020)).
 Therefore, we analysed the potential impact of choosing a suboptimal lambda with a decrease in evaluation count.
-The breakdown includes a subset of languages on both tasks (using the XLM-R version).
-We chose the zero-shot scenario where we performed language arithmetic between English and related language adapters.
-In Figure 3, we plot the validation scores with the corresponding baselines, that is, the scores of using directly the adapters.
+The breakdown includes a subset of languages on both tasks (using the XLM-R as a base model).
+We chose the zero-shot scenario where we performed LA between English and related language adapters.
+In Figures 5 and 8, we plot the validation scores with the corresponding baselines, that is, the scores of using directly the adapters.
 The dotted lines are based on $\lambda=0$ or $\lambda=1$ for clarity, meaning we exclusively use the arithmetic equation’s left or right side (i.e., a specific language).
@@
 At last, setting the default $\lambda=0.5$ would be near optimum for the analysed subset.
-Figure 3: Interpolation of $\lambda$ values for the zero-shot XLM-R scenario (NER - top, NLI - bottom; for QA see Appendix A) on the validation dataset.
+Figure 5: Interpolation of $\lambda$ values for the zero-shot XLM-R scenario (NER, for NLI and QA see Appendix D) on the validation dataset.
 The horizontal dashed lines represent the baseline scores for both languages used in language arithmetic.
 5.2 Language relatedness
+$LA_{l_{1}}(l_{1}\downarrow,l_{2}\rightarrow)$ ca es fr it pt ro
+Eval language NER
+es 73.82 - 75.06 74.61 74.90 73.68
+fr 75.74 75.76 - 75.82 75.59 75,79
+[…]
+es 78.23 - 78.04 77.96 77.94 78,02
+fr 77.95 77,65 - 77.70 77,47 77,60
+Table 1: Impact of language relatedness on the language arithmetic.
+We compare different Romance languages as a right side of $LA$ equation, i.e.
+$l2$ (both tasks use XLM-R model).
+We report an average over three runs.
 Relatedness of languages is a difficult-to-define concept.
@@
 Afterwards, we evaluated languages shared in our NER and NLI tasks (Spanish and French) by arithmetic with the entire Romance languages pool.
-The results are presented in Table 3 and show that given a related language (in this case, defined as coming from the same language family), there might be very little difference in which one we will use.
-The relative difference between the best and the worst language reaches only a $0.40$ F1 score in the NER task and the same value in terms of accuracy points in NLI.
-Nevertheless, the more sophisticated and hand-crafted language choice would improve the results from Table 2.
-Table 3: Impact of language relatedness on the language arithmetic. We compare different Romance languages as a right side of $LA$ equation, i.e. $l2$ (both tasks use XLM-R model).
-$LA(l_{1}\downarrow,l_{2}\rightarrow)$ ca es fr it pr ro
-Eval language NER
-es 78.77 - 78.96 79.06 78.91 78.66
-fr 76.63 76.92 - 76.70 76.80 76.80
-[…]
-es 77.07 - 77.23 77.13 77.07 76.83
-fr 76.77 76.99 - 76.97 76.95 77.07
-5.3 Language vs task vectors
-Task vectors exhibit high sparsity and orthogonality, as Ilharco et al.
-[19] observed.
-While the former characteristic can be denoted in language vectors (Figure 4), the latter displays different properties, in contrast to task vectors.
-In Figure 5, we visualise the cosine similarity between evaluated language vectors of language adapters.
-Notably, the minimal cosine similarity ($0.2$) surpasses the maximum ($0.18$) reported by previous research in the task space [19].
-Interestingly, most pairs in the task space oscillate within the range of $0.01$ to $0.03$.
-At the same time, language vectors surpass $0.2$ in each case, indicating that the orthogonality aspect is an inherent property of task adapters.
-Figure 4: Language vectors, similar to task vectors, are extremely sparse.
-The kernel density estimate plot presents the weights of a Spanish mBERT adapter.
-The behaviour is consistent across sampled layers and languages.
-Based on the cosine similarity observation, we investigated one of the recent task arithmetic extensions, Ties-Merging [44].
-This work introduces a three-step algorithm that prevents different parameter interferences, improving upon task arithmetic.
-The algorithm decreases the cosine similarity via a pruning step and alignment of parameter signs to perform arithmetic only on relevant parameters to the merged tasks.
-On the experimental details note, as Ties-Merging operates on averaging, not addition, we utilise a different lambda range during validation (as suggested by Yadav et al.
-[44], $\lambda\in[0.8,1.8]$), and we set Top-Kto the default value of $20$ and additionally to $80$.
-Figure 5: Cosine similarity between language vectors of language adapters.
-We report the comparison in the zero-shot setting on the NER task (XLM-R version) in Table 4.
-The Ties-Merging decreases the results significantly; it is not only outperformed by language arithmetic in every language pair but also by other baselines (see Table 1).
-Moreover, we note that the pruning operation has the reverse effect; higher pruning (i.e.
-keeping Top-Klower) decreases the performance (in contrast to task vectors) by making language vectors more sparse and, hence, closer to orthogonal.
-Table 4: Ties-Merging evaluation in the zero-shot setup on the NER task (XLM-R version).
-In the case of language arithmetic, where the language vectors have a higher overlap (i.e.
-higher cosine similarity), removing parameter interference decreases the overall performance.
-Method AVG F1 score
-LA 61.20
-Ties-Merging (Top-K20) 56.40
-Ties-Merging (Top-K80) 59.88
-One interpretation of the phenomena can be the different goals of the arithmetic: in the multi-task setup, we try to include multiple, often disconnected, tasks into a single task vector.
-In contrast, the language vectors’ goal is to include the knowledge of the closely related language rather than remove the harmful artefacts.
-Our findings indicate that language arithmetic has different characteristics than task arithmetic, and the follow-up works that improve upon task arithmetic might not be suited for the multilingual context.
+The results are presented in Table 1 and show that given a different related language (in this case, defined as coming from the same language family), there are minor scores fluctuation.
+The relative difference between the best and the worst language reaches around $1$ F1 score in the NER task and around $0.3$ in terms of accuracy points for the NLI task.
+This experiment indicates that a more sophisticated or hand-crafted language choice would improve the downstream results presented in Figure 5.
+However, it also shows that there is no free lunch, and results depend on a downstream task.
+For example, for Spanish evaluation, a French adapter is trained on the largest out of the listed languages raw corpora; therefore, for the NER task, it can leverage a bigger pool of seen during training Named Entities (at times language-independent or similar across languages) and perform the best given more data, even when there are closer related languages (according to methodology from Section 4.1.1 and Indo-European languages family tree).
 6 Related Work
@@
 However, unlike our work, the solutions require substantial changes to the vanilla adapter setup.
-The previous work either requires additional parameters to learn a parameterised composition function/a gating module to combine/steer the flow through the suitable adapter(s), or needs a specific training procedure that increases the complexity of the overall solution or, in most cases, both [33, 30, 22, 26, 9, 20, 41].
-Moreover, to prevent specifically negative interference, hyper-adapters [7] were proposed using hyper-networks [14], and Ansell et al.
-[3] applied sparse fine-tuning to compose task and language masks.
+The previous work either requires additional parameters to learn a parameterised composition function/a gating module to combine/steer the flow through the suitable adapter(s), or needs a specific training procedure that increases the complexity of the overall solution or, in most cases, both Pham et al.
+(2020); Pfeiffer et al.
+(2021); Lee et al.
+(2022); Parović et al.
+(2022); Chronopoulou et al.
+(2023b); Klimaszewski et al.
+(2023); Wang et al.
+(2023).
+Moreover, to prevent specifically negative interference, hyper-adapters Baziotis et al.
+(2022) were proposed using hyper-networks Ha et al.
+(2017), and Ansell et al.
+(2022) applied sparse fine-tuning to compose task and language masks.
 Unlike the prior studies mentioned earlier, our attempt is training-free and does not modify the base architecture.
 The most conceptually similar work is proposed by Chronopoulou et al.
-[8]; however, they operate on the notion of sample similarity to a subset of domains in a domain adaptation regime.
+(2023a); however, they operate on the notion of sample similarity to a subset of domains in a domain adaptation regime.
 Additionally, concurrent to our work, Parović et al.
-[27] show intial potential of task arithemtic in cross-lingual transfer based on a full fine-tuning setup.
+(2024) show intial potential of task arithemtic in cross-lingual transfer based on a full fine-tuning setup.
+However, in our work we focus on PEFT methods with additional, in-depth analysis.
 At last, we denote the rise of task arithmetic use cases, e.g.
-vision tasks or cross-task generalisation [39, 18].
+vision tasks or cross-task generalisation Stoica et al.
+(2023); Huang et al.
+(2024).
 7 Conclusion
 We have proposed language arithmetic, which enhances language adapters based on task arithmetic learning via addition.
-It is a training-free method and functions as a post-processing technique for the adapters.
-Our experiments have shown that LA is particularly beneficial in zero-shot and low-resourced cases.
+It is a training-free method and functions as a post-processing technique for MAD-X.
+Our experiments have shown that LA is particularly beneficial in a zero-shot scenario, where we do not have access to a target language adapter.
 At last, we highlight the differences between language and task arithmetic.
-In our future work, we plan to extend language arithmetic by incorporating additional components into the sum.
-Additionally, we aim to adapt other elements of the task arithmetic framework, initially designed for tasks, i.e.
+In our future work, we plan to extend language arithmetic by incorporating more components into the sum.
+Additionally, we aim to adapt other elements of the task arithmetic framework, i.e.
 task analogies and forgetting via negation, to a multilingual setup with an analysis of the differences between multi-task and multilingual arithmetic context.
@@
 Moreover, applying multi-source training based on the work of Ansell et al.
-[4] could provide better robustness of the task adapters and a more thorough analysis.
-At last, our analysis of the differences between task and language vectors is preliminary, without definitive conclusions.
-Appendix A Lambda impact - QA
-Figure 6 presents the analysis of lambda impact for the Question Answering task. For details, refer to Section 5.1.
-Figure 6: Interpolation of $\lambda$ values for the zero-shot QA XLM-R scenario on the validation dataset.
+(2023) could provide better robustness of the task adapters and a more thorough analysis.
+Appendix A Related languages
+We present the list of related languages used in our experiments in Table 2 (details in Section 4.1.1).
+Lang. ar bg de el es fr hi
+Related sw ru fr es fr es ur
+[…]
+Related bg ar bg hi ru ar
+Table 2: Languages used in the experiments with corresponding related languages. Details are provided in Section 4.1.1.
+Appendix B Zero-shot evaluation
+Figure 6: Zero-shot mBERT language arithmetic evaluation, where one side of the arithmetic is an English adapter, and the other is related to the target language adapter (e.g.
+French for Spanish - $LA_{es}(en,fr)$).
+The values above bars present a relative difference to a better proxy.
+Figure 6 presents the results of the experiments described in Section 4.2 for the mBERT model.
+Appendix C Improving existing language adapters
+Figure 7 presents the results of the experiments described in Section 4.3 for the mBERT model.
+Figure 7: Variants of language arithmetic compared to the MAD-X method in the use-case to improve an existing target language adapter.
+The values above bars present a difference between a better LA setup and the MAD-X framework for the mBERT model.
+Appendix D Lambda impact - NLI and QA
+Figure 8 presents the analysis of lambda impact for NLI and QA tasks. For details, refer to Section 5.1.
+Figure 8: Interpolation of $\lambda$ values for the zero-shot NLI and QA XLM-R scenario on the validation dataset.
 The horizontal dashed lines represent the baseline scores for both languages used in language arithmetic.
-Appendix B Related languages choice
-We derived related languages for the language adapters based on the available pool, i.e.
-we reused the adapters from other languages from the task (we trained two bonus ones, Persian and Italian, apart from the analysis in Section 5.2).
-While the related language could have been more suitable in a few cases, there is always a trade-off between relatedness, criterion for relatedness and available corpora.
-Moreover, the perfectly related language matching would be an unfair comparison as it would assume the availability of the ideal adapter’s pool, which is unlikely in practice.
-Table 5 presents the related language choice used in our experiments.
-Table 5: Languages used in the experiments with corresponding related languages.
-In the most cases, we used a related language from the already existing pool (i.e.
-evaluated languages, the first rows labelled “Lang.”); however, we trained some additional ones, including Italian and Persian.
-Lang. ar bg de el es fr hi
-Related Lang. fa ru fr de it it ur
-[…]
-Related Lang. bg ar de hi zh vi
+Appendix E Language vs task vectors
+Task vectors exhibit high sparsity and orthogonality, as Ilharco et al.
+(2023) observed.
+While the former characteristic can be denoted in language vectors (Figure 9), the latter displays different properties, in contrast to task vectors.
+In Figure 10, we visualise the cosine similarity between evaluated language vectors of language adapters.
+Notably, the minimal cosine similarity ($0.19$) surpasses the maximum ($0.18$) reported by previous research in the task space Ilharco et al.
+(2023).
+Interestingly, most pairs in the task space oscillate within the range of $0.01$ to $0.03$.
+At the same time, language vectors surpass $0.2$ in almost each case, indicating that the orthogonality aspect is an inherent property of task adapters.
+Figure 9: Language vectors, similar to task vectors, are extremely sparse.
+The kernel density estimate plot presents the weights of a Spanish mBERT adapter.
+The behaviour is consistent across sampled layers and languages.
+Based on the cosine similarity observation, we investigated one of the recent task arithmetic extensions, Ties-Merging Yadav et al.
+(2023).
+This work introduces a three-step algorithm that prevents different parameter interferences, improving upon task arithmetic.
+The algorithm decreases the cosine similarity via a pruning step and alignment of parameter signs to perform arithmetic only on relevant parameters to the merged tasks.
+On the experimental details note, as Ties-Merging operates on averaging, not addition, we utilise a different lambda range during validation (as suggested by Yadav et al.
+(2023), $\lambda\in[0.8,1.8]$), and we set Top-Kto the default value of $20$ and additionally to $80$.
+Figure 10: Cosine similarity between language vectors of language adapters.
+We report the comparison in the zero-shot setting on the NER task (XLM-R version) in Table 3.
+The Ties-Merging decreases the results significantly compared to the default language arithmetic.
+Moreover, we note that the pruning operation has the reverse effect; higher pruning (i.e.
+keeping Top-Klower) decreases the performance (in contrast to task vectors) by making language vectors more sparse and, hence, closer to orthogonal.
+Method AVG F1 score
+LA 60.54
+Ties-Merging (Top-K20) 52.94
+Ties-Merging (Top-K80) 57.57
+Table 3: Ties-Merging evaluation in the zero-shot setup on the NER task (XLM-R version, averaged over three runs and all evaluated languages).
+In the case of language arithmetic, where the language vectors have a higher overlap (i.e.
+higher cosine similarity), removing parameter interference decreases the overall performance.
+One interpretation of the phenomena can be the different goals of the arithmetic: in the multi-task setup, we try to include multiple, often disconnected, tasks into a single task vector.
+In contrast, the language vectors’ goal is to include the knowledge of the closely related language rather than remove the harmful artefacts.
+Our findings indicate that language arithmetic has different characteristics than task arithmetic, and the follow-up works that improve upon task arithmetic might not be suited for the multilingual context.